OpenStudy (anonymous):

√5 - 3 ------- √5 - 2

OpenStudy (anonymous):

how could you show 1/2 as a different fraction? same concept here

OpenStudy (anonymous):

multiply both numerator and denominator by \[\sqrt{5} + 2\]

OpenStudy (anonymous):

am i wrong?

OpenStudy (anonymous):

then rationalise and u'll get the ans say if u need working

OpenStudy (anonymous):

basketmath, u r on the wrong track in this question we have to rationalise the denominator

OpenStudy (anonymous):

right... so you would multiply both. that's where i was going with it

OpenStudy (anonymous):

if it were \[\sqrt[3]{5} - 2\] you would need to multiply by sqrt(5)^2+2 right?

OpenStudy (anonymous):

in yr number above is it sqrt5 or cubert5????

OpenStudy (anonymous):

cubert of 5

OpenStudy (anonymous):

@marguerite final answer is -(root5 +1) or -root5 -1

OpenStudy (anonymous):

to get 5 out of the cubert you would need to multiply by sqrt(5)^2

OpenStudy (anonymous):

so how would i write that?

OpenStudy (anonymous):

the final answer or ???

OpenStudy (anonymous):

the final answer

OpenStudy (anonymous):

\[-(\sqrt{5} +1)\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

can you show your work i got: \[\sqrt{5}^2 - 3 \over 2\]

OpenStudy (anonymous):

-2*

OpenStudy (anonymous):

@basketmath for yr previous query we will have to multiply it by cuberoot(5^2) + 2

OpenStudy (anonymous):

ok ill show the working

OpenStudy (anonymous):

root5 - 3 root5 + 2 --------- x --------- root5 - 2 root5 + 2 (root5)^2 + 2root5 - 3root5 -6 ---------------------------- (root5)^2 - 2^2 5 - root5 - 6 ----------- 5 - 4 -root5 - 1 --------- 1 -(root5 +1)

OpenStudy (anonymous):

Hope it is clear now.....:)

OpenStudy (anonymous):

Hey marguerite, if it is clear for u, pls click the good answer button for me

OpenStudy (anonymous):

Hey basketmaths, where r u???? hv u seen the working????